OpenStudy (anonymous):

how do you find the slope of the line passing through a pair of points? example (-15, 10) and (20, -5) ??

OpenStudy (anonymous):

from -15 to 20 is right 35 units from 10 to -5 is down 15 units slope is \(-\frac{15}{35}=-\frac{3}{7}\)

OpenStudy (lgbasallote):

or if you'd like a formula... \[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\] where m = slope \(y_2\) means second y coordinate \(y_1\) means first coordinate \(x_2\) means second x coordinate \(x_1\) means first x coordinate

OpenStudy (anonymous):

thanks so much, really helped!